OpenStudy (anonymous):

Laura paddles in her canoe at a speed of 5 mph for 5 miles. She then hops on her windsurfer and sails for 12 miles at a speed of 6 mph. For how long does Laura travel?

OpenStudy (anonymous):

She is traveling at 5mph for 5 miles and 5/5 = 1 so she travels for one hour in her canoe. Then she is traveling at 6mph for 12 miles and 12/6 = 2 so she travels for 2 hours on her windsurfer. 1+2 = 3 So Laura travels for 3 hours in total.